What mass of solution would be needed to deliver 3.1 g of a drug if the concentration of the drug in the solution was 3.5%.
1
Expert's answer
2020-11-09T11:52:35-0500

3.5% concentration means 3.5 g of solute in 100g of solution

3.5/100= 3.1/x

x= 100×3.1/3.5= 88.57g